OpenStudy (anonymous):

Use the quadratic formula to solve the equation x^2-3x-7=0

OpenStudy (anonymous):

Well the quadratic formula is \[x = \frac{ -b \pm \sqrt{b ^{2} - 4\left( a \right)} \left( c \right) }{ 2a }\] a would be the number that x^2 is being multiplied by (so in this case a=1), b is the number infront of x (so b=-3) and c is your last number (which is -7). Substitute and figure it out